OpenStudy (anonymous):

Tiara wants to divide a square piece of paper into two equalivalent triangles. If the square measures 20 cm on each side. What will the third side of each triangle measure?

jimthompson5910 (jim_thompson5910):

Hint: draw out a picture to get  Solve for x to get your answer

OpenStudy (anonymous):

so would it be 40 cm?

OpenStudy (anonymous):

Well you add the 90 degree angle

jimthompson5910 (jim_thompson5910):

You can use the pythagorean theorem to solve for x a^2 + b^2 = c^2 20^2 + 20^2 = x^2 ...

OpenStudy (anonymous):

would be 40^2

OpenStudy (anonymous):

then you use the right angle formula of \[a^2 + b ^2 = c^2 = c = \sqrt{400 + 400}\]

OpenStudy (anonymous):

\[c = \sqrt{800}\]

OpenStudy (anonymous):

so would it be 8.9 then?

jimthompson5910 (jim_thompson5910):

square root of 800 and not 80

OpenStudy (anonymous):

^

OpenStudy (anonymous):

28.3

jimthompson5910 (jim_thompson5910):

if you have to round to the nearest tenth, then yes

OpenStudy (anonymous):

well it shouldnt be 20 should it??

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

nvm Thank you